OpenStudy (anonymous):

a^2+b^2=25,and a^3+b^3=91.find a and b

OpenStudy (anonymous):

see ur solution is following

OpenStudy (anonymous):

do u know what is the formula of a^3 + b^3

OpenStudy (anonymous):

do u mean (a+b)(a+b)(a+b)=(a^2=2ab+b^2)(a+b)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then i am lost

OpenStudy (anonymous):

THE VALUE OF a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (anonymous):

a^2 + b^2 = 25

OpenStudy (anonymous):

(a+b)(25-ab)

OpenStudy (anonymous):

91 = (a+b)(25-ab)

OpenStudy (anonymous):

are u understanding

OpenStudy (anonymous):

yap following

OpenStudy (anonymous):

u must know that 3^2 + 4^2

OpenStudy (anonymous):

= 25

OpenStudy (anonymous):

put the value of a and b respectively to 3 and 4 in the equation a^3 + b^3 = 91

OpenStudy (anonymous):

u will get 27 + 64 that is 91 so a = 3 and b=4

OpenStudy (anonymous):

did u understand

OpenStudy (anonymous):

if yes than give me medal for this solution thanx

OpenStudy (anonymous):

yap.tq sir

OpenStudy (anonymous):

give me medal

OpenStudy (anonymous):

hey do u knw me if no than how u said me sir

OpenStudy (anonymous):

i use to it by calling my teacher "sir".how to give you medal

OpenStudy (anonymous):

click on the icon GIVE MEDAL on the right side of the message i m giving u and also need not say me sir i m in 8th grade and u???????????

OpenStudy (anonymous):

bye